Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 304     Entry time: 13 Aug 2023, 21:33
Author: Jayanath Koliyadu 
Group: SPB 
Subject: Shift end 

X-ray delivery

  • 9.3 keV, up to 181 pulses at 564 kHz, 8 - 1.3 mJ
  • Issue with pulse energy stability along the pulse train in the morning but optimzed later on in the day

Optical laser delivery

  • PP Laser, 800nm, 50 fs, up to 181 pulses @ from 10 Hz to564 kHz

Achievements/Observations

  • Optimized Bismuth diffraciton condition on AGIPD
  • Collected pump-probe data with Bismuth fixed target
  • Aligned PAM and measured arrival time with electron bunch arrival time slope enabled
    • Positive slope and V shape slope measured

Issues

  • Issues with beam stability and intenisty.
    • Optimized later on in the day
    • But did not recover what we had yesterday in terms of pulse energy
ELOG V3.1.4-7c3fd00